# Canvil: Designerly Adaptation for LLM-Powered User Experiences

### v1

@@ Abstract.
Advancements in large language models (LLMs) are poised to spark a proliferation of LLM-powered user experiences. In product teams, designers are often tasked with crafting user experiences that align with user needs. To involve designers and leverage their user-centered perspectives to create effective and responsible LLM-powered products, we introduce the practice of designerly adaptation for engaging with LLMs as an adaptable design material. We first identify key characteristics of designerly adaptation through a formative study with designers experienced in designing for LLM-powered products ($N=12$). These characteristics are to 1) have a low technical barrier to entry, 2) leverage designers’ unique perspectives bridging users and technology, and 3) encourage model tinkering. Based on this characterization, we build Canvil [^1] Canvil is available for public use at https://www.figma.com/community/widget/1277396720888327660., a Figma widget that operationalizes designerly adaptation. Canvil supports structured authoring of system prompts to adapt LLM behavior, testing of adapted models on diverse user inputs, and integration of model outputs into interface designs. We use Canvil as a technology probe in a group-based design study (6 groups, $N=17$) to investigate the implications of integrating designerly adaptation into design workflows. We find that designers are able to iteratively tinker with different adaptation approaches and reason about interface affordances to enhance end-user interaction with LLMs. Furthermore, designers identified promising collaborative workflows for designerly adaptation. Our work opens new avenues for collaborative processes and tools that foreground designers’ user-centered expertise in the crafting and deployment of LLM-powered user experiences.
@@
copyright: acmcopyrightjournalyear: 2024doi: xxxxxxx.xxxxxxxconference: arXiv; January; 2024price: 15.00isbn: xx-x-xxx-xxxx-xxxxxccs: Human-centered computing Interaction design process and methodsccs: Human-centered computing Empirical studies in collaborative and social computing
@@ 1. Introduction
A paradigm shift is underway for integrating artificial intelligence (AI) capabilities into everyday user-facing technologies. Large pre-trained AI models, most notably large language models (LLMs), have versatile natural language capabilities that unlock novel interactive techniques and interfaces for more intuitive and customizable user experiences across a wide spectrum of applications (Suh et al., 2023; Angert et al., 2023; Wu et al., 2022b; Parnin et al., 2023). However, these promises also come with numerous concerns. Integrating LLMs into a domain without careful consideration of the user contexts surrounding model use and implementation of behavioral guardrails for the model may result in user experiences that perpetuate societal biases (Santurkar et al., 2023), threaten users’ sense of well-being (Shah, 2023; Roose, 2023), or otherwise do harm (Selbst et al., 2019; Do et al., 2023).
@@
As technology development practitioners, designers [^2] We consider designers to be anyone who has an active, hands-on role in designing the user experience of a product and/or feature. This includes job titles such as UI and UX designer/researcher, content designer, and product designer. are uniquely positioned to mitigate these concerns (Feng and McDonald, 2023; Subramonyam et al., 2021a; Liao et al., 2023; Yang, 2018; Yildirim et al., 2022). Designers’ work often involves aligning technological capabilities with user needs, such that the technology addresses (or makes progress towards addressing) pain points identified in user research (Feng and McDonald, 2023; Subramonyam et al., 2022). Designers are trained in human-centered design methods that allow them to understand users and usage contexts, prototype potential solutions with relevant technology, and iterate on those solutions based on their understanding of users or user feedback (Ball, 2005). Yet, prior work has shown that designers face diverse challenges working with AI (Yang et al., 2020; Subramonyam et al., 2021b). These challenges include procedural ones, such as difficulties collaborating with the engineering teams training the models (Subramonyam et al., 2022, 2021b), as well as instrumental ones, such as lacking means to effectively work with the models (Feng and McDonald, 2023; Yang, 2018; Subramonyam et al., 2021a). In efforts to address this, researchers have situated AI as a design material to highlight exploring material considerations—e.g., the technology’s capabilities, limitations, and adaptability (Leonardi, 2012)—for designers to better understand and apply AI in the context of their design problems (Yang, 2018; Dove et al., 2017; Feng et al., 2023b). A good “designerly understanding” of AI can help designers ideate on new AI-powered design ideas, mitigate AI’s varying impact for different user scenarios, collaborate with design and non-design practitioners, and reinforce user-centered perspectives amongst the team and throughout the product development cycle (Liao et al., 2023; Wang et al., 2023).
@@
The advent of LLMs introduces new opportunities for engaging with AI as a design material. First, adaptability emerges as a key materialistic property of LLMs. LLMs are responsive to adaptation via fine-tuning (Dodge et al., 2020; Hu et al., 2021; Ouyang et al., 2022b; Bai et al., 2022) and prompt-based methods (OpenAI, 2023c; Wu et al., 2022a; OpenAI, 2023b). In fact, due to resource-intensive training of LLMs, it has become a standard practice for individual developers and development teams to adapt “base” LLMs from a small handful of providers (e.g., Anthropic, Google, OpenAI) (Parnin et al., 2023) for improved performance in domain-specific tasks. Moreover, natural language interaction and adaptation democratizes AI experimentation for practitioners—including designers—traditionally excluded from AI conversations due to limitations in technical expertise (Liao et al., 2023; Petridis et al., 2023b). Despite this, there has been limited exploration of designers’ interaction with—let alone adaptation of—LLMs in practice, with untapped opportunities for designers to contribute to LLM-powered product development.
In this paper, we introduce the practice of designerly adaptation of LLMs (henceforth “designerly adaptation”) to showcase new opportunities for designers to harness LLMs as an adaptable design material. We first characterize this new practice through a formative interview study with 12 designers experienced in designing for LLM-powered products and features. Our study reveals that desirable characteristics of designerly adaptation include execution via natural language prompt-based methods, leveraging designers’ user-centered perspectives and expertise, and encouraging iterative tinkering with models.
While these desiderata align with those surfaced in prior work (Feng and McDonald, 2023; Subramonyam et al., 2022; Yang et al., 2020), the practical significance of satisfying them in design workflows remains unclear. To investigate, we build Canvil, a technology probe in the form of a Figma [^3] Figma is a popular collaborative design tool: https://www.figma.com/. widget that operationalizes designerly adaptation. Canvil enables designers to adapt LLMs within their Figma canvases via system prompting (Microsoft, 2023; OpenAI, 2023c) and integrate outputs from adapted models into their designs. We use Canvil in a task-based design study with 6 groups of 17 designers total to understand whether and how designerly adaptation can assist in crafting LLM-powered user experiences. We find that through designerly adaptation, designers steered LLM behavior with user needs, derived interface affordances to enhance user interaction with LLMs, and recognized promises in collaborative adaptation to share resources and knowledge with design and non-design stakeholders. They were optimistic about integrating designerly adaptation into their own workflows and noted procedural and ethical questions to address in practice.
Equipped with findings from both studies, we propose a concrete workflow for designerly adaptation to offer a tangible launchpad for researchers and practitioners to further explore, critique, and iterate on this practice. We end by underscoring the potential of tools for collaborative AI tinkering, while reflecting on materiality’s impact on social practices within product teams.
Concretely, our work makes the following contributions:
@@
  • Insights from a formative study with designers experienced in crafting LLM-powered user experiences, from which we assemble a characterization of designerly adaptation—a new practice by which designers engage with LLMs as an adaptable design material.
@@
  • Insights from an empirical exploration of how designers engage in designerly adaptation in a task-based design study.
  • A discussion of our work’s implications on (collaborative) design and beyond, including a proposed workflow for designerly adaptation to orient future work.
@@ 2. Related Work
To situate our work, we review prior literature in AI as a design material, approaches to LLM adaptation, and tools for interactively working with AI models.
@@ 2.1. AI as a Design Material
Robles and Wiberg argue that recent advances in computational technologies bring about a “material turn”—a transformation within interaction design that allows for the shared use of material metaphors (e.g., flexibility) across physical and digital worlds (Robles and Wiberg, 2010). Indeed, prior works have discussed AI as a design material (Dove et al., 2017; Yang, 2018; Yang et al., 2020; Benjamin et al., 2021; Luciani et al., 2018; Liao et al., 2023; Feng and McDonald, 2023; Feng et al., 2023c; Yildirim et al., 2022), and highlighted why AI’s materiality can make it uniquely difficult to design with (Yang et al., 2020; Subramonyam et al., 2021b; Liao et al., 2023; Dove et al., 2017; Benjamin et al., 2021). AI is often treated as a black box to non-technical stakeholders such as designers, making it challenging to tune user interactions to often unpredictable and complex model behavior (Benjamin et al., 2021; Yang, 2018). In addition, AI’s technical abstractions are often divorced from concepts designers are familiar with (Subramonyam et al., 2021b), and designers consequently struggle with creatively using or manipulating the material to generate design solutions (Liao et al., 2023; Feng and McDonald, 2023). AI models are also non-deterministic and fluid in nature—they may evolve with new data or user input, and can be intentionally steered towards desirable behaviors with choices of data, algorithm, parameter, and so on (Yang et al., 2020). Without a concrete material understanding to begin with, designers are unable to grasp the nature of these uncertainties (Yang, 2018; Subramonyam et al., 2021a) or the opportunities to shape the design materials for desirable UX (Liao et al., 2023). Yang et al. (Yang et al., 2020) showed that these challenges of working with AI as a design material persist through the entire double-diamond design process—from identifying the right user problem to be solved by AI to designing the right UX to solve the user problem.
@@
Recent advances in LLMs simultaneously alleviate and exacerbate some of the aforementioned challenges. The barrier to tinkering with AI has significantly lowered thanks to the use of natural language as a primary mode of interaction and easily accessible tools such as ChatGPT. Concerns, however, have also arisen over the lack of explainability and transparency in LLMs due to their complex technical architectures (Liao and Vaughan, 2023). Yet, because of LLMs’ powerful capabilities, there is significant interest in exploring their integration into user-facing technologies (Pichai, 2023; Stallbaumer, 2023; Parnin et al., 2023); as such, designers should be prepared to work with them as a design material (Kulkarni et al., 2023). Despite this, we have yet to understand designers’ current approaches and desiderata when crafting LLM-powered user experiences. Our work contributes both empirical knowledge and tooling via a technology probe to this space.
@@ 2.2. Adaptation of Large Language Models
Adaptation can also occur without modifying the model itself, through system prompting. [^4] System prompting has also been referred to as “system messaging” and “metaprompting” (Microsoft, 2023). We use “system prompting” in our paper to align with terminology in prior academic work (Wei et al., 2023). Different from one-off prompting, system prompting applies to all individual user inputs for how the model should behave (e.g., “always respond in a concise manner”), often targeting an application domain. OpenAI’s API provides a “System” field to specify system prompts (OpenAI, 2023a, c), which have been shown to be effective at modifying model behaviors (OpenAI, 2023b; Salewski et al., 2023; Deshpande et al., 2023). For instance, a line of work explored instructing the model to behave with a certain “persona” to elicit desirable or adversarial model behaviors (Wei et al., 2023; Cheng et al., 2023; Salewski et al., 2023; Wolf et al., 2023; Deshpande et al., 2023). However, system prompts can be challenging to author (Zamfirescu-Pereira et al., 2023), and there is not yet an established “gold standard” for prompt writing, system or otherwise. Researchers and practitioners have attempted to derive useful prompting formats based on empirical exploration; these include in-context learning (i.e., by providing desired input-output examples) (Brown et al., 2020; Wu et al., 2023; Zhao et al., 2021), chain-of-thought reasoning (Wei et al., 2022), and instruction-following (Kaddour et al., 2023; Ouyang et al., 2022a). Industry recommendations have also emerged for system prompts, encouraging the specification of elements such as context specification (e.g., “You are Yoda from Star Wars”), task definition (e.g., “You respond to every user input as Yoda and assume the user is a Padawan”), and safety guardrails (e.g., “If the user requests inappropriate or offensive responses, you must respectfully decline with a wise Yoda saying”) (Microsoft, 2023).
In our work, we examine and support designerly adaptation of LLMs—whether and how designers can contribute to model adaptation for user-facing products through system prompting. We develop a characterization of the term through our formative study, and then build a technology probe to operationalize it in design practice. In building our probe, we draw from recommended best practices for system prompting, but acknowledge that these practices may shift over time and that our work reflects just one possible instantiation of these practices.
@@ 2.3. Interactive Tools for Steering AI Behavior
Literature at the intersection of HCI and AI has introduced a wide range of interactive techniques and tools to aid humans in training and adapting AI models, ranging from ones supporting data scientists to perform data wrangling (Görtler et al., 2022; Wang et al., 2022; Kandel et al., 2011; Liu et al., 2020; Wongsuphasawat et al., 2015), model training and evaluation (Bäuerle et al., 2022; Cabrera et al., 2023; Amershi et al., 2015; Ren et al., 2016; Microsoft, 2024; Google, 2024; IBM, 2024), managing model iterations (Hohman et al., 2020; Robertson et al., 2023), and so on, to those allowing for “human-in-the-loop” paradigms at various stages of the model development pipeline, including data annotation (Smith et al., 2018; Settles, 2012; Chao et al., 2010; Monarch, 2021; Raghavan et al., 2006), output correction (Branson et al., 2010; Bodén et al., 2021), integration testing (Chen et al., 2022), and explainability (Lage et al., 2018; Qian et al., 2019; Nguyen and Choo, 2021).
@@
There has also been a long-standing interest in “democratizing AI” for domain experts or practitioners without formal technical training to steer model behaviors. Interactive machine learning (iML) (Fails and Olsen, 2003) is a field responding to this interest by advocating for interactive and incremental model steering through intuitive interfaces and tightly coupled input-evaluation feedback loops. For instance, interfaces for transfer learning (Mishra and Rzeszotarski, 2021) have enabled non-expert users to transfer learned representations from a larger model to a domain-specific task. Tools have also encouraged non-expert exploratory tinkering of AI models through visual drag-and-drop UIs (Carney et al., 2020; Liner.ai, 2022; Lobe.ai, 2021; Du et al., 2023). Teachable Machine (Carney et al., 2020) is one such tool that allows users to train models for image, video, and audio classification. Rapsai (Du et al., 2023) is a visual programming pipeline for rapidly prototyping AI-powered multimedia experiences such as video editors. In an era where formal knowledge about AI is limited to a small handful of technical experts, these tools also serve to demystify AI for everyday users.
@@
With the onset of language models, barriers to experimenting with AI have lowered even further, paving the path for a new generation of interactive AI tools. Sandbox environments such as OpenAI’s playground (OpenAI, 2024) require no prerequisites besides a grasp of natural language to prompt the model and steer model behavior. However, relying on unstructured natural language alone can be daunting and ineffective. Many tools have risen to the challenge to support prompt engineering with more structured tinkering. Prompt chaining is one such approach (Wu et al., 2022b, a; Arawjo et al., 2023; Suh et al., 2023), by which users can use a node-based visual editor to write prompts for simple subtasks and assemble them to solve a larger, more complex task. PromptMaker (Jiang et al., 2022) and MakerSuite (Huffman and Woodward, 2023) allow the user to rapidly explore variable-infused prompts and few-shot prompting, while ScatterShot (Wu et al., 2023) helps specifically with curating few-shot prompting examples.

### v2

@@ Abstract.
Advancements in large language models (LLMs) are sparking a proliferation of LLM-powered user experiences (UX). In product teams, designers often craft UX to meet user needs, but it is unclear how they engage with LLMs as a novel design material. Through a formative study with 12 designers, we find that designers seek a translational mechanism that enables design requirements to shape and be shaped by LLM behavior, motivating a need for designerly adaptation to facilitate this translation. We then built Canvil [^1] Canvil is available for public use at https://www.figma.com/community/widget/1277396720888327660., a Figma widget that operationalizes designerly adaptation. We used Canvil as a technology probe in a group-based design study (6 groups, $N=17$), finding that designers constructively iterated on both adaptation approaches and interface designs to enhance end-user interaction with LLMs. Furthermore, designers identified promising collaborative workflows for designerly adaptation. Our work opens new avenues for processes and tools that foreground designers’ user-centered expertise in LLM-powered applications.
@@
copyright: acmcopyrightjournalyear: 2024doi: xxxxxxx.xxxxxxxconference: arXiv; September; 2024price: 15.00isbn: xx-x-xxx-xxxx-xxxxxccs: Human-centered computing Interaction design process and methodsccs: Human-centered computing Empirical studies in collaborative and social computing
@@ 1. Introduction
A paradigm shift is underway for integrating artificial intelligence (AI) capabilities into everyday user-facing technologies. Large pre-trained AI models, most notably large language models (LLMs), have versatile natural language capabilities that unlock novel interactive techniques and interfaces for more intuitive and customizable user experiences across a wide spectrum of applications (Suh et al., 2023; Angert et al., 2023; Wu et al., 2022b; Parnin et al., 2023). However, these promises also come with numerous concerns. Integrating LLMs into a domain without careful consideration of the user contexts surrounding model use and implementation of appropriate guardrails may result in user experiences that perpetuate societal biases (Santurkar et al., 2023), threaten users’ sense of well-being (Shah, 2023; Roose, 2023), or otherwise do harm (Selbst et al., 2019; Do et al., 2023; Weidinger et al., 2022; Chan et al., 2023; Shelby et al., 2023).
@@
As technology development practitioners, designers [^2] We consider designers to be anyone who has an active, hands-on role in designing the user experience of a product and/or feature. This includes job titles such as UX designer/researcher, content designer, and product designer. are uniquely positioned to mitigate these concerns (Feng and McDonald, 2023; Subramonyam et al., 2021a; Liao et al., 2023; Yang, 2018; Yildirim et al., 2022). Designers’ work often involves aligning technological capabilities with user needs, such that the technology addresses (or makes progress towards addressing) pain points identified in user research (Feng and McDonald, 2023; Subramonyam et al., 2022). Designers are trained in human-centered design methods that allow them to understand users and usage contexts, prototype potential solutions with relevant technology, and iterate on those solutions based on their understanding of users or user feedback (Ball, 2005). Yet, prior work has shown that designers face diverse procedural challenges (e.g., difficulties collaborating with the engineering teams training the models (Subramonyam et al., 2022, 2021b)) and instrumental challenges (e.g, lacking means to effectively work with the models (Feng and McDonald, 2023; Yang, 2018; Subramonyam et al., 2021a)) when working with AI (Yang et al., 2020; Subramonyam et al., 2021b). In efforts to address these challenges, researchers have situated AI as a design material to highlight its material properties—including the technology’s capabilities, limitations, and adaptability (Leonardi, 2012)—for designers to better understand and apply AI in the context of their design problems (Yang, 2018; Dove et al., 2017; Feng et al., 2023b). A good designerly understanding of AI can help designers ideate on new AI-powered design ideas, mitigate AI’s varying impact for different user scenarios, collaborate with design and non-design practitioners, and reinforce user-centered perspectives amongst the team and throughout the product development cycle (Liao et al., 2023; Wang et al., 2023).
@@
The advent of powerful pre-trained LLMs introduces new opportunities for engaging with AI as a design material. First, adaptability emerges as a key materialistic property of LLMs—these models are responsive to adaptation via fine-tuning (Dodge et al., 2020; Hu et al., 2021; Ouyang et al., 2022b; Bai et al., 2022) and prompt-based methods (OpenAI, 2023c; Wu et al., 2022a; OpenAI, 2023b). In fact, due to resource-intensive training of LLMs, it has become a standard practice for individual developers and development teams to adapt “base” LLMs from a small handful of providers (e.g., Anthropic, Google, OpenAI) instead of training custom in-house models (Parnin et al., 2023). Moreover, natural language interaction and adaptation democratizes AI experimentation for practitioners—including designers—traditionally excluded from AI conversations due to limitations in technical expertise (Liao et al., 2023; Petridis et al., 2023b). Despite this, there has been limited exploration of designers’ interaction with—let alone adaptation of—LLMs in practice as they contribute to LLM-powered features and products.
In this paper, we first conduct a formative interview study with 12 designers experienced in designing features and products that use LLMs to better understand their workflows and desiderata when crafting LLM-powered UX. From our interviews, we identify a need for a translational mechanism that allow design requirements to directly shape model behavior, and vice versa. Drawing from Cross’ notion of design codes—processes or frameworks that connect design knowledge (often embedded within design materials) with design requirements (Cross, 1982)—we propose designerly adaptation for LLMs (henceforth “designerly adaptation”) as a new design code to facilitate this two-way translation. We operationalize designerly adaptation by building Canvil, a Figma [^3] Figma is a popular collaborative design tool: https://www.figma.com/. widget that supports structured authoring of model behavioral specifications based on design requirements to adapt LLMs, testing of adapted models on diverse user inputs, and integration of model outputs into interface designs. To investigate whether and how designerly adaptation can assist in crafting LLM-powered UX in practice, we used Canvil as a technology probe in a task-based design study with 6 groups of 17 designers. We find that through using Canvil, designers iteratively translated diverse user and design requirements into LLM behavioral specifications. This allowed them to simultaneously adapt their designs to enhance user interaction with LLMs, while setting or refining behavioral constraints for the model using their designs. Furthermore, designers recognize promises in collaboratively engaging with designerly adaptation to share resources and knowledge with design and non-design stakeholders. They are optimistic about integrating designerly adaptation into their own workflows and note additional procedural and instrumental questions to address in practice. Concretely, our work makes the following contributions:
@@
  • Insights from a formative study with designers experienced in crafting LLM-powered user experiences, from which we motivate designerly adaptation—a design code bridges material exploration of LLMs with design processes.
@@
  • Insights about how designers engage in designerly adaptation from an empirical exploration of designers using Canvil in a task-based design study.
  • A discussion of our work’s implications on design practice and beyond, including a proposed workflow for designerly adaptation to orient future work and reflection of LLMs’ sociomaterial implications on collaborative practices within product teams.
@@ 2.1. AI as a Design Material
Robles and Wiberg argue that recent advances in computational technologies bring about a “material turn”—a transformation within interaction design that allows for the shared use of material metaphors (e.g., flexibility) across physical and digital worlds (Robles and Wiberg, 2010). Indeed, prior works have discussed AI as a design material (Dove et al., 2017; Yang, 2018; Yang et al., 2020; Benjamin et al., 2021; Luciani et al., 2018; Liao et al., 2023; Feng and McDonald, 2023; Feng et al., 2023c; Yildirim et al., 2022), and highlighted why AI’s materiality can make it uniquely difficult to design with (Yang et al., 2020; Subramonyam et al., 2021b; Liao et al., 2023; Dove et al., 2017; Benjamin et al., 2021). AI is often treated as a black box to non-technical stakeholders such as some designers, making it challenging to tune user interactions to often unpredictable and complex model behavior (Benjamin et al., 2021; Yang, 2018). In addition, AI’s technical abstractions are often divorced from concepts designers are familiar with (Subramonyam et al., 2021b), and designers consequently struggle with creatively using or manipulating the material to generate design solutions (Liao et al., 2023; Feng and McDonald, 2023). AI models are also non-deterministic and fluid in nature—they may evolve with new data or user input, and can be intentionally steered towards desirable behaviors with choices of data, algorithms, parameters, and so on (Yang et al., 2020). Without a concrete material understanding to begin with, designers face challenges grasping the nature of these uncertainties (Yang, 2018; Subramonyam et al., 2021a) or the opportunities to shape the design materials for desirable UX (Liao et al., 2023). Yang et al. (Yang et al., 2020) showed that these challenges of working with AI as a design material persist through the entire double-diamond design process—from identifying the right user problem to be solved by AI to designing the right UX to solve the user problem.
@@
Recent advances in LLMs simultaneously alleviate and exacerbate some of the aforementioned challenges. The barrier to tinkering with AI has significantly lowered thanks to the use of natural language as a primary mode of interaction and easily accessible tools such as ChatGPT. Concerns, however, have also arisen over the lack of transparency and controllability in LLMs due to their complex technical architectures (Liao and Vaughan, 2023). Yet, because of LLMs’ powerful capabilities, there is significant interest in exploring their integration into user-facing technologies (Pichai, 2023; Stallbaumer, 2023; Parnin et al., 2023); as such, designers should be prepared to work with them as a design material (Kulkarni et al., 2023). Despite this, we have yet to understand designers’ current approaches and desiderata when crafting LLM-powered user experiences. Our work contributes both empirical knowledge and tooling via a technology probe to this space.
@@ 2.2. Adaptation of Large Language Models
Adaptation can also occur without modifying the model itself via system prompting (Microsoft, 2023; Wei et al., 2023; OpenAI, 2023b; Salewski et al., 2023; Deshpande et al., 2023). Different from one-off user prompting, system prompting applies to all individual user inputs for how the model should behave (e.g., “always respond in a concise manner”), often targeting an application domain. For instance, a line of work explored instructing the model to behave with a certain “persona” to elicit desirable or adversarial model behaviors (Wei et al., 2023; Cheng et al., 2023; Salewski et al., 2023; Wolf et al., 2023; Deshpande et al., 2023). However, system prompts can be challenging to author (Zamfirescu-Pereira et al., 2023), and there is not yet an established “gold standard” for prompt authoring, system or otherwise. Researchers and practitioners have attempted to derive useful prompting formats based on empirical exploration; these include in-context learning (i.e., by providing desired input-output examples) (Brown et al., 2020; Wu et al., 2023; Zhao et al., 2021), chain-of-thought reasoning (Wei et al., 2022), and instruction-following (Kaddour et al., 2023; Ouyang et al., 2022a). Industry recommendations have also emerged for system prompts, encouraging the specification of elements such as context specification (e.g., “You are Yoda from Star Wars”), task definition (e.g., “You respond to every user input as Yoda and assume the user is a Padawan”), and safety guardrails (e.g., “If the user requests inappropriate or offensive responses, you must respectfully decline with a wise Yoda saying”) (Microsoft, 2023).
In our work, we examine and support designerly adaptation of LLMs—a means of empowering designers to explore LLMs as a design material through model adaptation. While we draw from recent developments in and recommended best practices for accessible model adaptation via natural language, we acknowledge that these practices may shift over time and using currently available techniques is just one possible instantiation of designerly adaptation.
@@ 2.3. Interactive Tools for Steering AI Behavior
Literature at the intersection of HCI and AI has introduced a wide range of interactive techniques and tools to aid humans in training and adapting AI models, ranging from ones supporting data scientists to perform data wrangling (Görtler et al., 2022; Wang et al., 2022; Kandel et al., 2011; Liu et al., 2020; Wongsuphasawat et al., 2015), model training and evaluation (Bäuerle et al., 2022; Cabrera et al., 2023; Amershi et al., 2015; Ren et al., 2016; Microsoft, 2024; Google, 2024; IBM, 2024), managing model iterations (Hohman et al., 2020; Robertson et al., 2023), and so on, to those allowing for “human-in-the-loop” paradigms at various stages of the model development pipeline, including data annotation (Smith et al., 2018; Settles, 2012; Chao et al., 2010; Monarch, 2021; Raghavan et al., 2006), output correction (Branson et al., 2010; Bodén et al., 2021), and integration testing (Chen et al., 2022), and explainability.
@@
There has also been a long-standing interest in “democratizing AI” for domain experts or practitioners without formal technical training to steer model behaviors. Interactive machine learning (iML) (Fails and Olsen, 2003) is a field responding to this interest by advocating for interactive and incremental model steering through intuitive interfaces and tightly coupled input-evaluation feedback loops. For instance, interfaces for transfer learning (Mishra and Rzeszotarski, 2021) have enabled non-technical audiences to transfer learned representations in a model to a domain-specific task. Tools have also encouraged exploratory tinkering of AI models through visual drag-and-drop UIs (Carney et al., 2020; Liner.ai, 2022; Lobe.ai, 2021; Du et al., 2023). Teachable Machine (Carney et al., 2020) is one such tool that allows users to train models for image, video, and audio classification. Rapsai (Du et al., 2023) is a visual programming pipeline for rapidly prototyping AI-powered multimedia experiences such as video editors. In an era where formal knowledge about AI is limited to a small handful of technical experts, these tools also serve to demystify AI for everyday users.
@@
With the onset of language models, barriers to experimenting with AI have lowered even further, paving the path for a new generation of interactive AI tools. Sandbox environments such as OpenAI’s playground (OpenAI, 2024) require no prerequisites besides a grasp of natural language to prompt the model and steer model behavior. However, relying on unstructured natural language alone can be daunting and ineffective. Many tools have risen to the challenge to support prompt engineering with more structured tinkering. Prompt chaining is one such approach (Wu et al., 2022b, a; Arawjo et al., 2024; Suh et al., 2023), by which users can use a node-based visual editor to write prompts for simple subtasks and assemble them to solve a larger, more complex task. PromptMaker (Jiang et al., 2022) and MakerSuite (Huffman and Woodward, 2023) allow the user to rapidly explore variable-infused prompts and few-shot prompting, while ScatterShot (Wu et al., 2023) helps specifically with curating few-shot prompting examples.